OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! Solve the inequality:

OpenStudy (anonymous):

\[-1+4y \le 31\]

OpenStudy (michele_laino):

hint: if I add +1 to both sides, I get this: \[\Large - 1 + 4y + 1 \leqslant 31 + 1\] please simplify

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

\[4y \le 32\]

OpenStudy (anonymous):

what do i do after that

OpenStudy (michele_laino):

ok! Now if I divide both sides by 4, I get this: \[\Large \frac{{4y}}{4} \leqslant \frac{{32}}{4}\] again, please simplify

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got \[y \le8\]

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thank you for your help!

OpenStudy (michele_laino):

:)